OpenStudy (catscatscats):

The period of function f(x) = sec(x/4) is

OpenStudy (catscatscats):

2Π radians Π radians 4Π radians 8Π radians

OpenStudy (amistre64):

what is the period normally associated with the sec function?

OpenStudy (catscatscats):

2pi

OpenStudy (amistre64):

then all we need to do is solve for x when: x/4 = 2pi

OpenStudy (catscatscats):

ummm

OpenStudy (amistre64):

you can ignore the pi part since all the options have it ... when is x/4 = 2 ?

OpenStudy (catscatscats):

the answer isn't 2pi

OpenStudy (amistre64):

when does x/4 = 2?

OpenStudy (catscatscats):

Idk? that's why im confused?

OpenStudy (amistre64):

if you are doing trig, then you must have graduated from algebra class ...

OpenStudy (amistre64):

something, divided by 4, is equal to 2 ...

OpenStudy (catscatscats):

Ugh yeah but like how does that have to do with the problem

OpenStudy (catscatscats):

The answer isn't 2pi. I've already tried it

OpenStudy (amistre64):

sec(2pi) is what we want to go for. sec(x/4) = sec(2pi), when x/4 = 2pi

OpenStudy (amistre64):

solve for x ...

OpenStudy (catscatscats):

so 4pi

OpenStudy (amistre64):

no, 4pi/4 is not 2pi, its pi

OpenStudy (catscatscats):

8pi.

OpenStudy (amistre64):

8pi/4 = 2pi ... yes

OpenStudy (catscatscats):

Thank you. Can you help with some more too please

OpenStudy (amistre64):

not at the moment, but good luck